two rods of lengths l1 and l2 are made of materials whose co-efficient of linear expansion are aan a2. ​If the increase in length be same at all temperatures then
(i) l1/l2=a1/a2
(ii) l1/l2=a2/a1
(iii)  l1 + a1 = l2 + a2 
(iv) l1/a12=l2/a22

Grade:11

1 Answers

arun
8 years ago
answer is (ii). it is given that change in length is same for both rods. now,
{\color{Red} \Delta l=laT}                           
here, red coloured l, a represents the l1 and a1
and green coloured represents l2 and a2 
{\color{Green} \Delta l=laT}
and it is given{\color{Green} \Delta l}={\color{Red} \Delta l}
\Rightarrow {\color{Red} lat}={\color{Green} lat}
also{\color{Red} t}={\color{Green}t}
\Rightarrow \frac{{\color{Red} l}}{{\color{Green} l}}=\frac{{\color{Green} a}}{{\color{Red} a}}
so l1/l2=a2/a1